OpenStudy (anonymous):

Let x(sub i) be a random variable uniformly distributed over the interval from zero to one, X(sub i)~U(0,1). Let the random variable y be related to X(sub i) as follows: Y Sigma (from i=1 to 108) X(sub i)

OpenStudy (anonymous):

\[\sum_{i=1}^{108}X _{i}\]

OpenStudy (anonymous):

What is the mean of Y? I have 54. it's the right answer

OpenStudy (anonymous):

What is the Variate of Y? i got 972.1666666666666

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

ah thank god

OpenStudy (jamesj):

The mean is definitely right. Write down your logic for Var \[ Var(Y) = E[(Y-E[Y])^2 \] = ...

OpenStudy (jamesj):

and forget the 108. Write that as n. The particular value is a distraction.

OpenStudy (anonymous):

basically just took the summation of (n-m)^2 / n

OpenStudy (anonymous):

or rather (ni -m)^2 / n where i is "sub i"

OpenStudy (anonymous):

\[\sum_{i=1}^{n}((n _{i}-\mu)^2)/2\]

OpenStudy (anonymous):

er /n

OpenStudy (zarkon):

If you are taking a random sample from some distribution with finite variance then \[Var(Y)=Var\left(\sum_{i=1}^{n}X_i\right)=\sum_{i=1}^{n}Var\left(X_i\right)\]